OpenStudy (anonymous):

the following is a graph of the parent function y= 1/x. T/F and explain to me. tia(:

OpenStudy (anonymous):

@jim_thompson5910 @loser66

jimthompson5910 (jim_thompson5910):

you need to post the graph

OpenStudy (xapproachesinfinity):

where the graph? lol

OpenStudy (anonymous):

it didnt post :/ there i think it should work this time

OpenStudy (anonymous):

@Loser66 @jim_thompson5910 @xapproachesinfinity

OpenStudy (anonymous):

@ganeshie8 @sammixboo @iambatman help por favor?

OpenStudy (sheraz12345):

Put in the values of x and get answers for y.. THEN plot em.. This is how the graph has been constructed

OpenStudy (anonymous):

Yes, that is the graph for y=1/x.

OpenStudy (anonymous):

but is that the same thing as the graph for the PARENT function y=1/x?

OpenStudy (anonymous):

@sheraz12345 sorry but that made no sense to me

OpenStudy (anonymous):

y=1/x is already the parent graph, so yes, it still is the same thing. :)

OpenStudy (anonymous):

okay. thank you all (: